OpenStudy (anonymous):

f(x)={ kx+1, x<1 and x^2, x is greater or equal to 1} a)find the value of k so that f(x) is continuous at x=1 b) using the value of k found part a) sketch the graph of f(x)

OpenStudy (anonymous):

in order for the function to be continuous, lim f(x) = 1 x->1^-

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

no:(

OpenStudy (anonymous):

this is much easier than it looks put \(x=1\) in both expressions, set them equal, solve for \(k\)

OpenStudy (anonymous):

so put k(1)+1=(1)^2 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I got k=0?

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

me 3

OpenStudy (anonymous):

how would I graph this?

OpenStudy (debbieg):

Just make a piecewise graph.... f(x)={ 1, x<1 ={x^2 x>=1 Both pretty basic functions. :)

OpenStudy (anonymous):

thank you:) i think i got it now